### v3

@@ 6.1 Comments on the POf and the DFPOm
For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 which ensure their validity. We did not find generic rules that ensure Assumptions 1 and 2, but nevertheless they are not very demanding in practice, as we discuss in Remarks 2 and 3. Moreover, they become tractable in some cases such as in Section 4. Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets. Moreover, a careful reading of the proof of Theorem 1 shows that Theorem 1.a) actually holds under no assumption except the applicability of the POf, and that Theorem 1.b) holds as soon as $\chi$ is continuous. Hence, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). This scenario arises in some nontrivial cases, for example from optimal control theory as in [10, Chapter 7.3.1]. We also remark that Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$) for all $x^{*}$. In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition. Sections 4.1 and 4.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.
@@ 6.2 Perspectives for future work
Two lines of improvements appear for extending the POf. The first concerns the oracle function $\gamma$, as it is unrealistic to assume that a global solution to each subproblem is computable. A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). This would be representative of most practical cases. We conducted some tests following this idea on some optimal control problems in [10, Chapter 7] by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) (chosen according to acknowledged guidelines from the dedicated literature). This approach seems promising and works well in our experiments, although the theory is technical so we are still working to simplify it. We may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{r}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{i}}$). This would allow tackling Reformulation ($\mathbf{P}_{\texttt{r}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
@@
Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{r}}$). Currently, Reformulation ($\mathbf{P}_{\texttt{r}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible. We observed in [10, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$. A progressive barrier [5] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{r}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). In all the problems we tested in [10, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
@@
Beyond these two main lines of improvement, we also plan to develop more advanced algorithms for the DFPOm. This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{r}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any $x\in X$). Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{i}}$) by optimizing jointly the partition set to consider and the point to consider into the set. We may also learn or dynamically adapt the partition and its indexation to improve the behaviour of $\Phi$ and the computational cost of $\gamma$.

### v5

@@ 7.1 Comments on the POf and the DFPOm
For theoretical considerations, our main contributions are Theorems 1 and 2, and Assumptions 1 and 2 ensuring their validity. Yet, the proof of Theorem 1 actually shows that our analysis is not tight. First, Theorem 1.a) holds under no assumption except the applicability of the POf, and Theorem 1.b) holds as soon as $\chi$ is continuous. Consequently, neither Assumptions 1 and 2 nor Algorithm 1 are required when we may compute analytically a (global or local) solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). This scenario arises in some nontrivial cases, as in Section 4. Second, we require that $\mathbb{Y}$ is a Banach space in the whole paper, but actually our analysis requires only for $\mathbb{Y}$ to be complete, and this is needed only for Theorems 1.c) and 2. As a consequence, it is even not mandatory to ensure the completeness of $\mathbb{Y}$ in Theorems 1.a) and 1.b). Hence, in Section 4, we could use any norm on $\mathcal{A}\mathcal{C}$. Third, Theorem 2 is stronger if we moreover assume that $\varphi$ is lower semicontinuous and $\Omega$ is closed. In this case, Theorem 2 ensures that $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$) for all $x^{*}$. In our analysis where this additional assumption does not hold, Theorem 2 only ensures that $\gamma(x^{*})$ is a generalized local solution, as the subsequent claim ensuring that $\gamma(x^{*})$ is a (usual) local solution relies on a sufficient but not necessary condition. Sections 5.1 and 5.3 illustrate cases where $\varphi$ is not semicontinuous and $\gamma(x^{*})$ is, respectively, a local solution nevertheless, or only a generalized local solution as claimed.
We did not find generic rules that ensure Assumptions 1 and 2. Yet, they are likely met in practice and not very demanding, as we discuss in Remarks 2 and 3. Moreover, Assumptions 1 and 2 become tractable in some cases, as in Section 5. Indeed, they hold if $\mathbb{X}$ admits a partition into ample continuity sets of both $\sigma$ and $\varepsilon$ (a light assumption similar to [3]), and $\widetilde{\varphi}$ is piecewise uniformly continuous and bounded below with bounded level sets. Finally, if the assumptions do not hold, the DFPOm becomes heuristic, but none of our experiments in [11, 10] exhibit a drop in the performance of the method.
Theorems 1 and 2 assume that $\gamma$ returns exact global solutions to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in\mathbb{X}$ and Algorithm 1 runs ad infinitum to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) exactly. None of these assumptions hold true in practice. Nevertheless, our numerical experiments show that the DFPOm remains efficient when $\gamma$ returns approximations of global solutions and Algorithm 1 is interrupted early. In Sections 5.4 and 6.4, $\gamma$ is defined as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$), for all $x\in X$, with high confidence to approximate a global solution. We also observe in [10] that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions. Then, such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions exactly and Algorithm 1 runs ad infinitum.
Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{ref}}$) using a DFO algorithm with a covering step. We consider the cDSM (Algorithm 2) in our numerical experiments, but any DFO algorithm may be enhanced with a covering step, as discussed in Section 3.2 and [3]. The choice of the most appropriate DFO algorithm is likely to depend on $\Phi$. In general, Reformulation ($\mathbf{P}_{\texttt{ref}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason to favour a DSM over other classes of methods. In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be the most suited method to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$). However, to our best knowledge, no popular DFO solver implements algorithms with a covering step yet. Then, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality), so they are not compatible with Theorem 2 yet. Nevertheless, [3] remarks that the covering step is likely optional in practice. We second this observation in Section 6.4, as solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) using Nomad provides better results than our naive cDSM, even though Nomad does not implement a covering step.
The usual case of application of the DFPOm is when $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. Moreover, $\dim(\mathbb{X})\leq 50$ is desirable when solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) with the cDSM, for performance reason [6, Section 1.4]. We stress that the performance of the DFPOm depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence, $\dim(\mathbb{Y})$ does not directly matter, so it may be possibly infinite as in [10]. Moreover, the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is, the larger the constant $\tau$ from Section 6 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$. We also remark that the way the partition is indexed influences the shape of $\Phi$. The solving process may be substantially simplified by a re-parametrization of $\mathbb{X}$ that avoid concentrations of discontinuities or mitigate local variations of $\Phi$.
@@ 7.2 Perspectives for future work
Two lines of improvements appear for extending the POf. The first concerns the oracle function $\gamma$. It is unrealistic to assume that a global solution to each subproblem is computable. A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). This would be representative of most practical cases. We test this idea in [11, Chapter 7] in a context of optimal control problems, by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method (chosen according to acknowledged guidelines from the literature) solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). This approach seems promising and works well in our experiments, although the theory is technical so we are still working to simplify it. We may also determine if, given $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$). This would allow tackling Reformulation ($\mathbf{P}_{\texttt{ref}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
@@
Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{ref}}$). Currently, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible. We observed in [11, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$. A progressive barrier [5] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). In all the problems we tested in [11, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
@@
Beyond these two main lines of improvement, we also plan to develop more advanced algorithms for the DFPOm. This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$). Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) by optimizing jointly the partition set to consider and the point to consider into the set. We may also learn or dynamically adapt the partition and its indexation to improve the behaviour of $\Phi$ and the computational cost of $\gamma$.